OpenStudy (anonymous):

Find the range if y = 3x + 12 for x = {0.4, 0.6, 1.8, 2.2}. {1.2, 1.8, 5.4, 6.6} {13.2, 13.8, 17.4, 18.6} {12.4, 12.6, 13.8, 14.2}

OpenStudy (anonymous):

K so for this problem, all you have to do is pluck in those x-values and solve. We'll do them one by one :) So when we put in x=0.4 we get, y=3(0.4)+12 (Now we do order of operations and multiply 3*0.4 first) y=1.2+12 and now we add this together and get y=13.2 Now we do the same thing with the other 3 and get: y=3(0.6)+12 y=1.8+12 y=13.8 Now 2 more left: y=3(1.8)+12 y=5.4+12 y=17.4 And the last one: y=3(2.2)+12 y=6.6+12 y=18.6 So by doing this we now know that all our y-values are 13.2,13.8,17.4,18.6 and that's the second answer :)

OpenStudy (anonymous):

ur teachin me stuff,u ready for round 3?

OpenStudy (anonymous):

Sure and hey you can just stay here if you want :) I just like helping people learn more about getting medals so it'll be easier for the both of us as well :)

OpenStudy (anonymous):

idk what this is trying to get me to do?? John began his job making $25 the first day. After that he was paid $6.75 per hour. Write the linear equation for this problem in y=mx+b form.

OpenStudy (anonymous):

Oh okay, so it wants us to put the information you wrote above in an equation that looks like that. So every hour we will get $6.75 so that number will always be constant. Since we don't know how many hours John works, that'll be a variable, x. So so far we have y=$6.75x. Now it also said he got an automatic $25 on the first day so the equation is y=$6.75x+25 :) Did you get it? :)

OpenStudy (anonymous):

oh x is how many hours and thats multiplied by how much we get n a hour??

OpenStudy (anonymous):

Well we can't solve the equation, we can only make one. To solve it, the problem would have had to say how many hours John actually worked.

OpenStudy (anonymous):

now its aking me to do stuff with the same prob but with a t graph ;( idk this stuff

OpenStudy (anonymous):

John began his job making $25 the first day. After that he was paid $6.75 per hour.Make a T-table for the equation for this, using the x-values: 0, 20, 40. Type in the ordered pairs you get as your answer, in this form: (0, number), (20, number), (40, number).

OpenStudy (anonymous):

He gets 6.75 an hours, and we don't know how many hours he worked so that's why we do y=6.75x+25. the X in the equation above is the amount of hours John worked. But we can't put a number for x because the problem doesn't tell us how many hours John worked.

OpenStudy (anonymous):

i get that but not this

OpenStudy (anonymous):

Oh I get it. You have to make a table. Now though you have to put 0,20, & 40 as the x-values and solve for your answer.

OpenStudy (anonymous):

So our equation we have is y=$6.75x+25. We will pluck in 0,20 & 40 one by one to see what we get. y=$6.75(0)+25 y=0+25 y=$25. y=$6.75(20)+25 y=135+25 y=$160 y=$6.75(40)+25 y=270+25 y=$295

OpenStudy (anonymous):

 I'm pretty sure that's what it's asking for.